OpenStudy (asdfghj8):

How do I solve: 9x^4 - 27x^6 Over 3x^3 ?????

jimthompson5910 (jim_thompson5910):

Hint: Factor out 3x^3 from the numerator \[\Large \frac{9x^4-27x^6}{3x^3}\] \[\Large \frac{3x^3*(3x-9x^3)}{3x^3}\] I'll let you finish up

OpenStudy (mathmate):

@asdfghj8 In math, we \(solve\) an equation, or \(simplify\) an expression. For example, we solve the following equation (two expressions separated by an equal sign). 4x=8 x=2 or we simply the following expression: 27x^2/(3x) =9x